OpenStudy (anonymous):

What percent of 240 is 60?

OpenStudy (anonymous):

star you have a good sense of humour :P

OpenStudy (katrinakaif):

What percent : Lets say that to be "n" Of = Multiplication Is = Equal n x 240 = 60

OpenStudy (katrinakaif):

Now simply solve =)

OpenStudy (anonymous):

is it 240/60

OpenStudy (katrinakaif):

Not quite. Try again

OpenStudy (anonymous):

i would say the opposite

OpenStudy (anonymous):

60/240

OpenStudy (anonymous):

you are dividing the 240 from n and whatever you do to one side you do to the other

OpenStudy (katrinakaif):

^Correct, Now simplify You can cancel out the "zeros" and simply have the expression to be 6/24 = 1/4

OpenStudy (anonymous):

you are trying to isolate the n

OpenStudy (anonymous):

then 1/4=0.25 thats 25%

OpenStudy (anonymous):

Bingo :)

hero (hero):

Pipa, are you related to Sopa?

hero (hero):

I couldn't resist

OpenStudy (anonymous):

Who is Sopa?

hero (hero):

lol nvm